OpenStudy (anonymous):

Let f(x) = 4x - 5 and g(x) = 6x - 3. Find F(x) - g(x). A. 2x + 5 B. -2x - 8 C. -2x - 2 D. 2x - 5

OpenStudy (anonymous):

@sidsiddhartha I got 10x-8 can you tell me what I did wrong ?

OpenStudy (anonymous):

I did 4x-5+6x-3

ganeshie8 (ganeshie8):

but you're asked to take the difference , not addition...

OpenStudy (anonymous):

Ohhh wait I was supposed to subtract duhhh !

OpenStudy (anonymous):

Lol yea I just realized that

OpenStudy (anonymous):

So it would be B right ?

ganeshie8 (ganeshie8):

hahah okay :) and no, try again

OpenStudy (anonymous):

Okay C ?

ganeshie8 (ganeshie8):

f(x) = 4x - 5 and g(x) = 6x - 3 F(x) - g(x) 4x-5 - (6x - 3)

ganeshie8 (ganeshie8):

Yep ! C is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

Thanks

ganeshie8 (ganeshie8):

yw

OpenStudy (anonymous):

Let f(x) = 3x + 2 and g(x) =x - 3. Find f(x) - g(x). A. 2x + 5 B. 2x - 5 C. 4x - 1 D. 2x - 1

OpenStudy (anonymous):

I got D is that right @ganeshie8